# Following the Human Thread in Social Navigation

### v1

@@ 4.1 Quantitative Results
4.2 Ablation studies
In the following section, we investigate the impact of various forms of privileged information and their adaptability. Additionally, we assess how the length of the trajectories fed in Stage 1 affects our model.
4.2.1 Adatptable information
Table 2 presents the results from Stage 1 (S1) and Stage 2 (S2) with the utilization of various privileged information, such as Humanoid GPS (hGPS) and human trajectories (traj.). During S1, particularly in $S$ and notably in $SPS$, incorporating hGPS as an additional input leads to superior results. This advantage is likely attributed to hGPS containing implicit information about human and robot positions, facilitating more efficient path selection, especially in $SPS$ scenarios.
However, this pattern is not evident in S2, where trajectories typically provide more adaptable information. As previously mentioned, hGPS inherently encompasses the robot’s position, posing challenges in regressing this data during the second stage due to the lack of initial context regarding the robot’s location relative to the environment. Furthermore, hGPS may be difficult to adapt when the human is detected and disappears around a wall. Since hGPS comprises polar coordinates, the distance between the robot and the human spans the wall. This issue does not arise when utilizing trajectories, as they do not require any information about the robot and can be adapted solely based on depth images and detection information.
Fig. 3(left) illustrates the trend of the First encounter step over episode during Stage 2. The plot confirms that using trajectories facilitates the robot to find the person, which it achieves after only approx. 450 steps, while the hGPS and hGPS+traj take more than 700 steps. Finding the robot first results in higher $S$ and $SPS$ metrics, but it may incur larger chances of collision ($CR$ metric), due to having to then follow it for longer, cf. Sec. 4.1.
@@ 4.2.1 Adatptable information
Timeframes S $\uparrow$ SPS $\uparrow$ F $\uparrow$ CR $\downarrow$ ES $\uparrow$
@@
4.2.2 Length of trajectories
Table 3 illustrates the influence of trajectory length on our SDA. Employing 5 past timeframes as input yields generally inferior results, possibly due to the insufficient historical data to capture human dynamics. Similarly, using 50 timeframes leads to comparable outcomes. Although $S$, $SPS$, and $F$ metrics improve relative to the 5-frame scenario, they still lag behind our proposed model. This could be attributed to potential redundancy and counterproductive encoding of information.
Refer to captionRefer to caption
Figure 5: We showcase two different episodes. On the left, the agent successfully follows the human after it swiftly moves in front of the door. On the right, an episode where the human decides to move backward and the agent steps back to make way.
4.3 Qualitative Results
We qualitatively demonstrate the results in our proposed SDA. Firstly, we showcase the agent’s ability to locate the humanoid within the environment by moving around, followed by its capability to follow the humanoid within the environment. Subsequently, we present two specific behaviors where the agent briefly spots the humanoid and one where it moves backward to create space for passage.
Fig. 4 shows an episode where the agent and the human are located in different rooms at the start. Then, the agent begins its search for the humanoid by navigating within the environment until the encounter takes place. After the encounter, the agent then transitions into the follow phase.
The episode in Fig. 5 (left) illustrates the agent’s capability to track the human even when it is briefly observed. As the agent searches for the target in the bathroom, the humanoid swiftly passes in front of the door (Step 145) and disappears from the agent’s view again (Step 150). Due to learned social dynamics, the agent exploits the humanoid’s behavior and begins to follow it.
In the right episode in Fig. 5, the agent has already located the human and its objective is to follow it. It is observed that at Step 340, the human decides to move backward, prompting the agent to move backward as well, creating the necessary space for passage by Step 360. Then, it continues its task of following.
Limitations.
Table 1 shows a dependency of stage 2 performance on the adaptability of information, highlighting the significance of this factor in determining the model’s effectiveness. Another limitation arises from having benchmarked the proposed model only in simulation and the reliance on information during training that is not easily accessible outside simulation (trajectories). Mitigating these limitations, the advent of Habitat 3.0 [36], has greatly minimized the domain gap between simulated and real environments.
5 Conclusions
Our study presents the Social Dynamics Adaptation model (SDA) for Social Navigation within interactive settings. Notably, it is the first to integrate privileged human dynamics information during training while adapting it in the following stage, enabling its application in realistic environments without relying on such privileged information. Our findings underscore the non-trivial nature of adapting information, highlighting the necessity for selective processes. In future research, we aim to extend our model by incorporating diverse human dynamics beyond trajectories, enhancing the robot’s comprehension of human movement patterns.

### v2

@@ A.8 Additional Analysis
Design Analysis.
RMA (Kumar et al., 2021) was used to adapt an agent from simulation to real-world deployment. In contrast, SDA encodes human behavior in trajectories, which can be inferred at test time to develop a socially aware navigation policy. While RMA encodes the environment configuration vector $e_{t}$ with information only at time $t$, we feed an entire trajectory encoding the position in the 20 timesteps up to time $t$. As shown in Table 10, encoding the human position only at time $t$ results in unsatisfactory performance, making the direct application of RMA relatively poor. In fact, it achieves just $3\%$ of Episode Success (ES) against the $43\%$ of SDA.
@@
# States Considered S $\uparrow$ SPS $\uparrow$ F $\uparrow$ CR $\downarrow$ ES $\uparrow$
1 0.55±0.01 0.14±0.02 0.03±0.01 0.75±0.01 0.03±0.01
@@
Given the sequential nature of the states that get encoded, we also evaluated Transformers and MLPs and ultimately selected the MLPs for SDA. In Table 11, we include this preliminary study. Note how the choice of the new sequence modeling mechanism is not trivial and strongly affects performance.
Table 11: Ablation on the type of encoder to consider during Stage 1.
Encoder Type S $\uparrow$ SPS $\uparrow$ F $\uparrow$ CR $\downarrow$ ES $\uparrow$
MLP (Proposed) 0.92±0.00 0.46±0.01 0.44±0.02 0.61±0.02 0.50±0.01
Transformer 0.85±0.01 0.15±0.02 0.27±0.01 0.76±0.01 0.12±0.01
A.9 Qualitative results
The episode in Fig. 7 (top) illustrates the agent’s capability to track the human even when it is briefly observed. As the agent searches for the target in the bathroom, the humanoid swiftly passes in front of the door (Step 145) and disappears from the agent’s view again (Step 150). Due to learned social dynamics, the agent exploits the humanoid’s behavior and begins to follow it. In the episode in Fig. 7 (bottom), the agent has already located the human and its objective is to follow it. It is observed that at Step 340, the human decides to move backward, prompting the agent to move backward as well, creating the necessary space for passage by Step 360. Then, it continues its task of following.
Figure 7: We showcase two different episodes. On the top, the agent successfully follows the human after it swiftly moves in front of the door. On the bottom, an episode where the human decides to move backward and the agent steps back to make way.